# Distributed Stochastic Gradient Descent with Staleness: A Stochastic Delay Differential Equation Based Framework

### v1

@@ IV-B Parameter Optimization for Pure Asynchronous SGD
If $N_{p}^{\kappa}(t)$ are mutually independent, [^14] Note that $N_{p}^{\kappa}(t)$ does not have to be a Poisson process. Also, by assuming that the computational delays of each worker are independent random variables, the arrival processes $N_{p}^{\kappa}(t)$ are also mutually independent renewal processes. $N_{p}(t)$ can approximated as a Poisson process, the rate of which is given by $\nu_{p}=\sum_{\kappa=1}^{K}\nu_{\kappa}$.
From Theorem 3, we directly have the following corollary.
Corollary 10
In pure asynchronous SGD, it holds that $\mathbb{E}\{\breve{\tau}\}=K-1$.
First, we aim to present the distribution of the gradient staleness and the optimal choice for the number of users in pure asynchronous SGD. Let $Y\sim\text{Pois}(\nu)$ be the random variable distributed according to a Poisson distribution with rate $\nu$. In this context, we have the following proposition.
@@
In pure asynchronous SGD, the discrete-time gradient staleness for Client $\kappa$’s gradient follows a mixed Poisson distribution, i.e.,
$\vspace{-1.1ex}\breve{\tau}_{\kappa}\sim\text{Pois}((\nu_{p}-\nu_{\kappa})T_{\kappa}).$ (30)
@@
Given $T_{\kappa}\sim\text{Exp}\left(\nu\right)$, the gradient staleness in asynchronous SGD with ideal communication channel follows a geometric distribution, i.e., $\breve{\tau}_{\kappa}\sim\text{Geo}\left(\frac{1}{K-1}\right),$ in which $K$ is the total number of workers. Further, we have $\mathbb{E}\{\breve{\tau}_{k}\}={K-1}$ and $\text{Var}\{\breve{\tau}_{k}\}=K(K-1)$.
@@
IV-C Optimal Worker Numbers in Asynchronous SGD Through an Ideal Data Link or Network
In this subsection, our objective is to determine the optimal choice of the number of workers in asynchronous SGD with a quadratic objective function $f(x)=\frac{1}{2}vx^{2}$, assuming an ideal data link. Here, an ideal data link refers to one that is error-free and zero-delay. In pure asynchronous SGD, the expectation of the step staleness is $\mathbb{E}\{\breve{\tau}\}=K-1$. Based on this fact, we present the following proposition.
@@ IV-C Optimal Worker Numbers in Asynchronous SGD Through an Ideal Data Link or Network
For pure asynchronous SGD through an ideal channel with a fixed learning rate, the optimal number of workers, denoted by $K^{*}$, is approximately given by
$\vspace{-1.5ex}K^{*}\approx 1+\frac{1}{v\eta e}.$ (31)
@@
Given the learning rate $\eta$, the two roots of the characteristics function can be determined by Proposition 2. Notice that it holds that $\text{Real}(\lambda_{0})<\text{Real}(\lambda_{1})$ when $\breve{\tau}\eta v<\frac{1}{e}$, and $\text{Real}(\lambda_{0})=\text{Real}(\lambda_{1})$ when $\breve{\tau}\eta v\geq\frac{1}{e}$. In this way, the convergence rate of the SGD algorithm is mainly determined by the dominant root $\text{Real}(\lambda_{0})$, i.e., $\mathbb{E}\{\hat{x}(t)\}\approx\hat{x}(0)e^{\lambda_{0}{K\nu t}},$ in which $\lambda_{0}=\frac{W_{0}{(-(K-1){v}\eta)}}{(K-1)\eta}.$ Such, $K^{*}=\arg\min_{K}\frac{K\text{Real}(W_{0}(-(K-1){v}\eta))}{(K-1)\eta}$. Since $\text{Real}\{W_{0}(s)\}$ achieves its minimum at $s=-\frac{1}{e}$, we arrive at the proposition. ∎
Remark:  In this discussion, we aim to explore the effect of the number of workers on the convergence rate with a worker number-aware step size in pure asynchronous SGD. Note that the expectation of the discrete-time gradient staleness is approximately given by $\mathbb{E}\{\breve{\tau}\}\approx K$ in which $K$ is the number of workers. In this case, we set $\eta=\eta_{0}/K$ for some constant $\eta_{0}$. The dominant characteristic root is given by $\lambda=W_{0}(-\eta_{0}v)/\eta_{0}$. Also, the update frequency is $\nu=\frac{K}{\mathbb{E}\{\hat{\tau}\}}$. Therefore, the elapsed time in SDDE is $\eta\nu t=\frac{\eta_{0}t}{\mathbb{E}\{\hat{\tau}\}}$. Therefore, the SGD evolves according $\mathbb{E}\{\hat{x}(t)\}\approx\hat{x}(0)\exp{\frac{W_{0}(-\eta_{0}v)t}{\mathbb{E}\{\hat{\tau}\}}}$. Therefore, it is seen that increasing $K$ does not effectively increase the performance in asynchronous SGD with a staleness-aware step size.
Also, given the learning rate $\eta$ and the total number of workers $K$, it can be seen that when $K\geq 1+\frac{1}{v\eta e}$ the performance of pure asynchronous SGD deteriorates due to the large step staleness. In this case, one could improve the performance by setting an appropriate $B$. A sensible choice of $B$ can be $B^{*}\approx K-\frac{1}{v\eta e}$ by Theorem 3.
IV-D Optimal Worker Number in Pure Asynchronous SGD Through a Bandwidth Constrained Channel Shared by Multiple Workers
In this subsection, we aim to determine the optimal number of workers in pure asynchronous SGD operating through a shared medium with limited bandwidth. In this case, the continuous-time gradient staleness is characterized by the sum of the computational delay and the communication delay. In this scenario, an increase in $K$ not only results in more frequent updates but also leads to a rise in the overall gradient delay due to increased communication delay.
@@
Further, the average communication delay, which is given by the sum of the average waiting time in the queue plus the average service time, is $\mathbb{E}\{D\}=\frac{1}{\mu_{Q}}\left(\frac{\rho}{2(1-\rho)}+1\right)$. In this way, the expectation of the continuous-time gradient staleness is given by the sum of the expectation of the computation and communication delay. More specifically, we have $\mathbb{E}\{\hat{\tau}\}=\frac{1}{\nu}+\frac{1}{\mu_{Q}}\left(\frac{\rho}{2(1-\rho)}+1\right).$ With limited bandwidth, the client can choose to start the computation of a new gradient right away after it finishes the computation of the previous one. It can also start the computation of the new gradient after it receives an ACK (acknowledgment) that the global model has been updated according to the gradient, which reduces the gradient staleness but also reduces the update frequency. In the latter situation, it still holds that $\mathbb{E}\{\breve{\tau}\}=K-1$. In the subsequent, we focus on the former situation and we have the following proposition.
@@
With $K$ workers in total, the expectation of step staleness in asynchronous SGD through a channel with a fixed service rate is approximately given by [^15] The arrival process of the gradients to the central server is the departure process of the $M/D/1$ queue. Rigorously, the departure process of the $M/D/1$ is not a Poisson process. But the intensity of the departure process equals the intensity of the arrival process to the queue as long as $\mu_{Q}>\nu$.
$\mathbb{E}\{\breve{\tau}_{\text{Q}}(K)\}\approx K\left(1+\frac{\nu}{2\mu_{Q}}\left(\frac{K\nu}{\mu_{Q}-K\nu}+2\right)\right),$ (32)
@@
Therefore, starting a computation of a new gradient without waiting for an ACK increases the average step staleness with a bandwidth limited channel. By Eq. (32), we present the following remark.
@@
Next, we have the following proposition on the optimal number of workers in pure asynchronous SGD,
@@
Given the learning rate $\eta$ and the queue capacity $\check{K}$, the optimal number of workers in pure asynchronous SGD can be determined as the smaller solution to the following quadratic equation:
@@
In practice, the workers can reduce the computation time by reducing the size of the minibatch, thus increasing the frequency of gradient updates. However, this approach can lead to significant bandwidth consumption and network congestion, consequently degrading algorithm performance. Moreover, a high-resolution quantization of the gradients with large gradient noise also leads to substantial bandwidth usage, thus being inefficient due to the waste of network bandwidth.
@@ IV-E Optimal Design for Event-Triggered Distributed Optimization
Given the state variable at $t_{0}$, denoted by $\boldsymbol{x}_{0}$, the determination of the triggering time can be formulated as the first exit time of a continuous-time stochastic process (4), i.e., $H_{E}=\inf\{t\geq t_{0}|\boldsymbol{x}(t)\in E(\boldsymbol{x}_{0}),\boldsymbol{x}(t_{0})=\boldsymbol{x}_{0}\},$ in which $E(\boldsymbol{x})=\{\tilde{\boldsymbol{x}}|\left\|\nabla f(\tilde{\boldsymbol{x}})-\nabla f(\boldsymbol{x})\right\|\geq\xi\}$. Following the discussion in Subsection III-B, we assume that the triggering time is a Gaussian random variable, i.e., $H_{E}\sim\mathcal{N}(\vartheta,\varsigma^{2})$. In this case, gradient staleness also arises in event-triggered SGD. Given $H_{E}=\zeta$, the conditional gradient staleness approximately follows a uniform distribution $U(0,\zeta)$. Therefore, the gradient staleness follows a compound distribution, i.e., $\tau\sim U(0,H_{E})$. Further, it is assumed that $\vartheta\gg\varsigma$. By the law of total expectation, it holds that $\mathbb{E}\{\tau\}=\frac{\vartheta}{2}$. Further, the p.d.f. of $\tau$ is given by $p_{\tau}(s)=\int_{s}^{\infty}\frac{1}{\sqrt{2\pi}\varsigma s}e^{-\frac{1}{2}\left(\frac{s-\vartheta}{\varsigma}\right)^{2}}{\rm{d}}s\text{, for }s\geq 0$. In this case, approximately uniformly distributed gradient staleness arises distributed SGD in with ETC when $\vartheta\gg\varsigma$. Next, for the solvable case, i.e., with the quadratic objective function $f(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}\boldsymbol{V}\boldsymbol{x}$, the expectation of the triggering-time in event-triggered SGD is approximately given by $\mathbb{E}\{H_{E}\}\approx\xi\sqrt{\frac{1}{\boldsymbol{x}_{0}^{T}\boldsymbol{V}^{4}\boldsymbol{x}_{0}}}$ when $\xi\ll\|\boldsymbol{x}_{0}\|_{2}$. Note that different from the previous discussion, SGD along different eigenvector directions can not be viewed as independent in event-triggered SGD.
@@
Remark:  Let $C$ be the total bandwidth allocated to the uplink channel in event-triggered SGD. More specifically, the workers can update up to $C$ gradients to the parameter server per second on average. We assume that the triggering threshold is carefully chosen such that the communication channel is fully utilized and the workers are assumed to be homogeneous. Recall that the expectation of the triggering-time is denoted by $\mathbb{E}\{H_{E}\}=\vartheta$. In this way, the gradient staleness in SDDE is approximately uniformly distributed, i.e., $\tau\sim U(0,\vartheta)$, in which $\vartheta=\frac{K\eta}{CT_{s}}$. Note that $\eta$ is the fixed learning rate and $T_{s}$ is the time slot length in distributed learning with event-triggered communication. Similar to the discussion in Subsection IV-D, the optimal number of workers is approximately the number of workers at which the gradient staleness minimizes the real part of the dominant root. In this way, we conclude that the optimal number of workers is approximately given by $K^{*}=\frac{C\varpi^{2}T_{s}}{\eta(e^{-\varpi}-1)}$ given a total bandwidth constraint in distributed learning with event-triggered communication. Given a total bandwidth limit, too many workers can induce a large degree of gradient staleness and the performance of event-triggered SGD can therefore deteriorate.
V Numerical Results
In this section, we shall present simulation and numerical results to verify the theoretical analysis for distributed SGD with stale gradient. We will further demonstrate the effect of gradient noise, step size, and gradient staleness on asynchronous SGD.
To illustrate the relationship between the roots of the characteristic function and the gradient staleness, Fig. 2(a) presents numerical results for the characteristic roots of Eq. (16) and Eq. (22). In particular, we consider a fixed gradient staleness, denoted by $\tau$. or a uniformly distributed gradient staleness, denoted by $\tau\sim\text{Uni}(0,\zeta)$, respectively. In both cases, the real part of the dominant root first decreases, then increases as $\tau$ or $\xi$ increases. With a fixed gradient staleness, the real part of the root achieves its minimum, i.e., $\text{Real}(\lambda_{0})=-e$ when $\tau=1/e$, while for uniformly distributed gradient staleness, the real part of the root achieves its minimum when $\zeta=0.6476$. The above results demonstrate that an SGD algorithm achieves performance gain with an appropriate gradient staleness. This may be counterintuitive that the gradient staleness does not always impede the convergence of the SGD method. However, the SGD process oscillates when $\tau>1/e$ or $\zeta$, and diverges when $\tau>\frac{\pi}{2}$ or $\zeta>\frac{\pi^{2}}{2}$. This illustrates that the convergence property considerably deteriorates with large a gradient staleness.

### v2

@@ IV-B Parameter Optimization for Pure ASGD
In pure ASGD, it holds that $\mathbb{E}\{\breve{\tau}\}=K-1$.
First, we aim to present the distribution of the gradient staleness and the optimal choice for the number of workers in pure ASGD. Let $Y\sim\text{Pois}(\nu)$ be the random variable distributed according to a Poisson distribution with rate $\nu$. In this context, we have the following proposition, which demonstrates that the discrete-time gradient staleness follows a mixed Poisson distribution [71].
@@
In pure ASGD, the discrete-time gradient staleness of Client $\kappa$’s gradient follows a mixed Poisson distribution, i.e.,
$\breve{\tau}_{\kappa}\sim\text{Pois}((\nu_{p}-\nu_{\kappa})T_{\kappa}).$ (49)
@@
Given $T_{\kappa}\sim\text{Exp}\left(\nu\right)$, the gradient staleness in ASGD with an ideal communication channel follows a geometric distribution, i.e., $\breve{\tau}_{\kappa}\sim\text{Geo}\left(\frac{1}{K-1}\right),$ in which $K$ is the total number of workers. Further, we have $\mathbb{E}\{\breve{\tau}_{k}\}={K-1}$ and $\text{Var}\{\breve{\tau}_{k}\}=K(K-1)$.
@@
IV-C Optimal Worker Numbers in ASGD Through an Ideal Data Link or Network
In this subsection, our objective is to determine the optimal choice of the number of workers in ASGD along the eigenvector with an eigenvalue $v>0$, assuming an ideal data link. Here, an ideal data link refers to one that is error-free and zero-delay. In pure ASGD, the expectation of the step staleness is $\mathbb{E}\{\breve{\tau}\}=K-1$. Based on this fact, we present the following proposition.
@@ IV-C Optimal Worker Numbers in ASGD Through an Ideal Data Link or Network
When $K\gg 1$, for pure ASGD through an ideal channel with a fixed learning rate, the optimal number of workers, defined as the number of workers under with the dominant root of SGD dynamics achieves its minimum, denoted by $K^{*}$, is given by
$K^{*}=1+\frac{1}{v\eta e}.$ (50)
@@
See Appendix D. ∎
Remark:  In this discussion, we aim to analyze the effect of the number of workers on the convergence rate of pure ASGD when using a worker number-aware step size. The expected value of the discrete-time gradient staleness is given by $\mathbb{E}\{\breve{\tau}\}=K-1$, in which $K$ denotes the number of workers. In this case, the worker number-aware step size is given by $\eta=\eta_{0}/K$, where the constant $\eta_{0}>0$. The dominant characteristic root is $\lambda_{0}=W_{0}(-\eta_{0}v)/\eta_{0}$. Further, the update frequency is $\nu={K\nu_{0}}$, where $\nu_{0}$ is the update frequency of one single worker. Consequently, by the dominant pole approximation, the SGD dynamics evolve as $\mathcal{O}(e^{{W_{0}(-\eta_{0}v)\nu_{0}t}})$ by the dominant pole approximation. This analysis indicates that increasing $K$ does not effectively improve performance in ASGD when using a worker number-aware step size.
Also, given the learning rate $\eta$ and the total number of workers $K$, it can be seen that when $K\geq 1+\frac{1}{v\eta e}$ the performance of pure ASGD deteriorates due to the large step staleness. In this case, one could improve the performance by setting an appropriate $B$. A sensible choice of $B$ can be $B^{*}=K-\frac{1}{v\eta e}$ by Theorem 2.
IV-D Optimal Worker Number in Pure ASGD Through a Bandwidth Constrained Channel Shared by Multiple Workers
In this subsection, we aim to determine the optimal number of workers in pure ASGD operating through a shared medium with limited bandwidth. In this case, the continuous-time gradient staleness is characterized by the sum of the computational delay and the communication delay. In this scenario, an increase in $K$ not only results in more frequent updates but also leads to a rise in the overall gradient delay due to increased communication delay.
@@ IV-D Optimal Worker Number in Pure ASGD Through a Bandwidth Constrained Channel Shared by Multiple Workers
Further, the average communication delay, which is given by the sum of the average waiting time in the queue plus the average service time, is $\mathbb{E}\{D\}=\frac{1}{\mu_{Q}}\left[\frac{\rho}{2(1-\rho)}+1\right]$. In this way, the expectation of the continuous-time gradient staleness is given by the sum of the expectation of the computation and communication delay. More specifically, we have $\mathbb{E}\{\hat{\tau}\}=\frac{1}{\nu}+\frac{1}{\mu_{Q}}\left[\frac{\rho}{2(1-\rho)}+1\right].$ With limited bandwidth, the client can choose to start the computation of a new gradient right away after it finishes the computation of the previous one. It can also start the computation of the new gradient after it receives an ACK (acknowledgment) that the global model has been updated according to the gradient, which reduces the gradient staleness but also reduces the update frequency. In the latter situation, it still holds that $\mathbb{E}\{\breve{\tau}\}=K-1$. In the subsequent, we focus on the former situation and we have the following proposition.
@@
With $K$ workers in total, the expectation of step staleness in ASGD through a channel with a fixed service rate is given by [^11] The arrival process of the gradients to the central server is the departure process of the $M/D/1$ queue. Rigorously, the departure process of the $M/D/1$ is not a Poisson process. But the intensity of the departure process equals the intensity of the arrival process to the queue as long as $\mu_{Q}>\nu$.
$\mathbb{E}\{\breve{\tau}_{\text{Q}}(K)\}=(K-1)\left[1+\frac{\nu}{2\mu_{Q}}\left(\frac{K\nu}{\mu_{Q}-K\nu}+2\right)\right],$ (51)
@@
Therefore, starting a computation of a new gradient without waiting for an ACK increases the average step staleness with a bandwidth-limited channel. By Eq. (51), we present the following remark.
@@
Next, we have the following proposition on the optimal number of workers in pure ASGD,
@@
Given the learning rate $\eta$ and the queue capacity $\check{K}$, the optimal number of workers in pure ASGD can be determined as the smaller solution to the following quadratic equation:
@@
In practice, the workers can reduce the computation time, i.e., the amount of time a worker takes to compute the gradient, by reducing the size of the minibatch, thus increasing the frequency of gradient updates. However, this approach can lead to significant bandwidth consumption and network congestion, consequently degrading algorithm performance. Moreover, a high-resolution quantization of the gradients with large gradient noise also leads to substantial bandwidth usage, thus being inefficient due to the waste of network bandwidth.
@@ IV-E Optimal Design for Event-Triggered Distributed Optimization
Given the state variable at $t_{0}$, denoted by $\boldsymbol{x}_{0}$, the determination of the triggering time can be formulated as the first exit time of a continuous-time stochastic process (5), i.e., $H_{E}=\inf\{t\geq t_{0}|\boldsymbol{x}(t)\in E(\boldsymbol{x}_{0}),\boldsymbol{x}(t_{0})=\boldsymbol{x}_{0}\},$ in which $E(\boldsymbol{x})=\{\tilde{\boldsymbol{x}}|\left\|\nabla f(\tilde{\boldsymbol{x}})-\nabla f(\boldsymbol{x})\right\|\geq\xi\}$. Following the discussion in Subsection III-B, we assume that the triggering time is a Gaussian random variable, i.e., $H_{E}\sim\mathcal{N}(\vartheta,\varsigma^{2})$. In this case, gradient staleness also arises in event-triggered SGD. Given $H_{E}=\zeta$, the conditional gradient staleness approximately follows a uniform distribution $U(0,\zeta)$. Therefore, the gradient staleness follows a compound distribution, i.e., $\tau\sim U(0,H_{E})$. Further, it is assumed that $\vartheta\gg\varsigma$. By the law of total expectation, it holds that $\mathbb{E}\{\tau\}=\frac{\vartheta}{2}$. Further, the p.d.f. of $\tau$ is given by $p_{\tau}(s)=\int_{s}^{\infty}\frac{1}{\sqrt{2\pi}\varsigma s}e^{-\frac{1}{2}\left(\frac{s-\vartheta}{\varsigma}\right)^{2}}{\rm{d}}s\text{, for }s\geq 0$. In this case, approximately uniformly distributed gradient staleness arises distributed SGD in with ETC when $\vartheta\gg\varsigma$. Next, for the solvable case, i.e., with the quadratic objective function $f(\boldsymbol{x})=\frac{1}{2}\boldsymbol{x}^{T}\boldsymbol{V}\boldsymbol{x}$, the expectation of the triggering-time in event-triggered SGD is given by $\mathbb{E}\{H_{E}\}=\xi\sqrt{\frac{1}{\boldsymbol{x}_{0}^{T}\boldsymbol{V}^{4}\boldsymbol{x}_{0}}}$ by assuming constant gradient $\boldsymbol{V}\boldsymbol{x}_{0}$ when $\xi\ll\|\boldsymbol{x}_{0}\|_{2}$. Note that different from the previous discussion, SGD along different eigenvector directions can not be viewed as independent in event-triggered SGD.
@@
Remark:  Let $C$ be the total bandwidth allocated to the uplink channel in event-triggered SGD. More specifically, the workers can update up to $C$ gradients to the parameter server per second on average. We assume that the triggering threshold is carefully chosen such that the communication channel is fully utilized and the workers are assumed to be homogeneous. Recall that the expectation of the triggering-time is denoted by $\mathbb{E}\{H_{E}\}=\vartheta$. In this way, the gradient staleness in SDDE is approximately uniformly distributed, i.e., $\tau\sim U(0,\vartheta)$, in which $\vartheta=\frac{K\eta}{CT_{s}}$. Note that $\eta$ is the fixed learning rate and $T_{s}$ is the time slot length in distributed learning with event-triggered communication. Similar to the discussion in Subsection IV-D, the optimal number of workers is approximately the number of workers at which the gradient staleness minimizes the real part of the dominant root. In this way, we conclude that the optimal number of workers is given by $K^{*}=\frac{C\varpi^{2}T_{s}}{\eta(e^{-\varpi}-1)}$ given a total bandwidth constraint in distributed learning with event-triggered communication. Given a total bandwidth limit, too many workers can induce a large degree of gradient staleness and the performance of event-triggered SGD can therefore deteriorate.
@@
V Numerical Results
In this section, we shall present simulation and numerical results to verify the theoretical analysis for distributed SGD with stale gradients. We will further demonstrate the effect of gradient noise, step size, and gradient staleness on ASGD.
To illustrate the relationship between the roots of the characteristic function and the gradient staleness, Fig. 2(a) presents numerical results for the characteristic roots of Eq. (26) and Eq. (41). In particular, we consider a fixed gradient staleness, denoted by $\tau$. or a uniformly distributed gradient staleness, denoted by $\tau\sim\text{Uni}(0,\zeta)$, respectively. In both cases, the real part of the dominant root first decreases, then increases as $\tau$ or $\xi$ increases. With a fixed gradient staleness, the real part of the root achieves its minimum, i.e., $\text{Re}\{\lambda_{0}\}=-e$ when $\tau=1/e$, while for uniformly distributed gradient staleness, the real part of the root achieves its minimum when $\zeta=0.6476$. The above results demonstrate that an SGD algorithm achieves performance gain with an appropriate gradient staleness. This may be counterintuitive that the gradient staleness does not always impede the convergence of the SGD method. However, the SGD process oscillates when $\tau>1/e$ or $\zeta$, and diverges when $\tau>\frac{\pi}{2}$ or $\zeta>\frac{\pi^{2}}{2}$. This illustrates that the convergence property considerably deteriorates with a large gradient staleness.